# Elemental Frequency-Based Supervised Classification Approach for the Search of Novel Topological Materials

### v1

@@ Abstract
The machine learning based approaches efficiently solves the goal of searching the best materials candidate for the targeted properties. The search for topological materials using traditional first-principles and symmetry-based methods often requires lots of computing power or is limited by the crystalline symmetries. In this study, we present frequency-based statistical descriptors for machine learning-driven topological material’s classification that is independent of crystallographic symmetry of wave functions. This approach predicts the topological nature of a material based on its chemical formula. With a balanced dataset of 3910 materials, we have achieved classification accuracies of 82% with the Support Vector Machine (SVM) model and 83% with the Random Forest (RF) model, where both models have trained on common frequency based features. We have checked the performances of the models using 5fold5-fold cross-validation approach. Further, we have validated the models on a dataset of unseen binary compounds and have efficiently identified 22 common materials using both the models. Next, we implemented the firstprinciplesfirst-principles approach to confirm the the topological nature of these predicted materials and found the topological signatures of Dirac, Weyl, and nodal-line semimetallic phases. Therefore, we have demonstrated that the implications of frequency-based descriptors is a practical and less complex way to find novel topological materials with certain physical post-processing filters. This approach lays the groundwork for scalable, data-driven topological property screening of complex materials.
@@ I INTRODUCTION
Materials discovery and design are essential in materials science, which involves finding particularly promising new materials, analyzing and optimizing their attributes for various applications. Topological materials (TMs)[1, 2, 3, 4, 5, 6, 7, 8, 9, 10] have become a subject of immense interest in recent years due to their robust edge/surface states protected by different symmetries from local perturbations such as disorder, impurities, magnetic flux, etc. [4, 11, 12]. Since the first realization of TMs in the form of the quantum spin Hall effect [13, 14], this field has remarkably expanded to cover a number of classifications, including topological insulators (TIs), topological semimetals (TSMs), etc. [15, 16, 17, 18, 19, 20, 21, 22] Since the inception of this new class of materials, a significant and persistent investigation has revolved around the methods for ascertaining whether a specific crystalline material possesses a unique topological signature or not. The method of addressing this question has primarily relied on foundational theoretical and experimental techniques that allow trial-and-error possibilities [5, 23, 24]. The first-principles computational approach and traditional experimental methods require a significant amount of time and resources [25, 26, 27, 28, 29, 30, 31]. Recently developed symmetry indicators and topological quantum chemistry theories allow the analysis of a large number of TMs with low computational requirements [32, 33, 34], but despite the effectiveness of these theories, they have limitations owing to the symmetries of the crystalline materials [3, 35]. For example, these theories failed to identify the topological nature of TaAs, which is a low-symmetry crystalline material and the first experimentally identified Weyl semimetal [23]. Moreover, the Chern insulators and the materials having Time Reversal Symmetries (TRS), etc.[14, 36, 37, 38] have robust non-trivial topological phases even when crystalline symmetries vanish. Therefore, we require novel approaches that overcome the aforementioned hurdles, highly accurate and easily executable for the exploration of TMs in condensed matter physics. The machine learning (ML) approach has become one of the adopted methods for screening materials with desired properties [39, 40]. This approach is computationally cheaper, data-driven, with high accuracy and precision, making it an indispensable tool in material discovery [34, 41, 42, 43, 44]. Various reports have been published in recent times that have predicted electronic and structural properties using regression and classification ML models based on chemical composition and crystal structures [35, 45, 46, 47, 48]. Recent advancements include graph-based methods such as crystal graph convolution neural networks (CGCNNs), which have achieved significant success in the prediction and identification of topological phases [49]. A machine-learned chemical descriptor has also been introduced by Ma $etal.$, thus contributing to the advancements and growing implementation of ML in TMs [50]. The frequency of an atom or an element in a given set of data is an important indicator of what type of elements are more feasible to form TMs. In this study, we use a frequency-based formulation to create a descriptor that can predict TMs with acceptable precision and accuracy. This formulation is straightforward, amenable to calculation, and understandable, and given materials can be categorized with an accuracy of more than 80% in topological trivial or non-trivial classes. This formulation is independent of the crystal symmetries of the materials and can be implemented on all types of materials. We have integrated our study using ML algorithms, $i.e.$, Support Vector Machine (SVM) and Random Forest (RF), followed by firstprinciplesfirst-principles calculations, for the sake of the verification of the predicted materials. This work aims to improve the accuracy and efficiency of ML models to classify materials for their non-trivial topological phases.
@@ II.1 Data preprocessing
These three quantities in Eq. (2) for each respective element are used to generate the three features for our ML models [refer to section 1.1 and Table S1 in supplementary information]. We generate these features by a weighted sum formula, giving three feature quantities $\mathit{f_{1}}$, $\mathit{f_{2}}$, and $\mathit{f_{3}}$ calculated from the quantities $\mathit{F_{x}^{nt}}$, $\mathit{F_{x}^{rt}}$, and $\mathit{\varDelta F}$ respectively.
@@ II.2 Machine Learning Implementation
Figure 1: The workflow of statistical frequency-based feature engineering approach for ML training and testing of models, and verification of predicted materials using first principles calculations.
@@ II.3 First-principles calculations
The electronic structure calculations based on the density functional theory (DFT) [53] were performed using the projector augmented-wave (PAW) [54] method as implemented in the Vienna ab initio Simulation Package (VASP) [55, 56, 57, 58, 59] code. The standard PAW potentials, preferably suggested in the VASP code, were used for all the elements for these calculations. The exchange and correlation energy were obtained with the generalized gradient approximation (GGA) of Perdew Burke Ernzerhof (PBE) functional [60], with the inclusion of SOC in the self-consistent field calculations [61]. The total energies convergence criterion of $\mathrm{10^{-6}}$ eV was adopted along with a fine gamma-centered k mesh. The plane-wave cut-off energies were set at 1.5 times the maximum energies available in the PAW potentials.
@@ III.1 Training and Testing
Figure 2: (a, b) Metrics of classification showing Confusion matrices and (c, d) Receiver Operating Characteristic (ROC) - Area Under the Curve (AUC) for the models SVM and RF, respectively.
@@ III.1.2 Cross Validation
Cross-validation is a key part of ML analysis, as it permits the removal of reliance on how the sample is divided into testing and training datasets for performance estimation. The cross-validation helps to discover and fix the problem of overfitting and underfitting, and it is a better way to estimate the generalization error of the training-testing split, as it tests the model on more than one validation dataset. This method is used to fine-tune the hyperparameters so they can pick the ones that work best for generalization. Cross-validation also helps make fair comparisons between different models by testing each one on several subsets of data under the same conditions. The complete dataset, consisting of $n$ elements, is randomly divided into $p$ subsets, each containing $n/p$ elements, referred to as folds, as shown in Table III. Here, we have carefully tested the performance of the SVM and RF models by using $\mathit{five-fold}$ cross-validation, and the AUC scores of each fold, as well as the mean AUC score, are shown in Table III.
@@ III.2 Analysis of data in $\mathrm{T_{1}}$
To support our frequency-based feature approach, we have analyzed the elemental data in the dataset $\mathrm{T_{1}}$ and calculated the elemental features $\mathit{f_{1}}$, $\mathit{f_{2}}$ using Eqs. (3) and (4), whereas the frequency difference $\mathit{\varDelta F}$ and feature $\mathit{f_{3}}$ have been calculated using Eqs. (2) and (5), respectively. We have observed that the $\mathit{\varDelta F}$ values are positive for the alkali metals group (with the exception of hydrogen), whereas the halogens have highly positive $\mathit{\varDelta F}$ values in the pblockp-block elements. Moreover, the highly electronegative elements, such as Oxygen, have an unusually high value of $\mathit{\varDelta F}$, which reflects their widespread occurrence across a varied range of materials in the dataset. The elementary distribution of the rest of the data for $\mathit{\varDelta F}$ across the periodic table can be seen in the heatmap shown in Fig. 3. Among the elements with high negative $\mathit{\varDelta F}$ values, Pd, Pt, Ti, Hf, Zr, U, Au, and Sc occur in more than 50% of the materials containing these elements and are labelled as topologically non-trivial. This observation and analysis imply a higher probability that materials containing these elements possess topological phases. Furthermore, the other elements with negative $\mathit{\varDelta F}$ values generally have a 30% to 50% chance of being in a material of a topological phase. Hence, it implies, statistically, that there is indeed a chance to form topological phases of materials when the elements with $\mathit{\varDelta F<0}$ are present in a material. Moreover, the elements with $\mathit{\varDelta F>0}$ have shown a less than 30% chance of contributing towards this phase. These conclusions are in line with the observation made in the dataset, where most elements with ionic characters are less likely to form topological phases due to their high electronegativity values. This reinforces the significance of the $\mathit{\varDelta F}$ metric as a discriminative feature in identifying potential topological materials and supports its integration into our ML framework. Table S2 in the supplementary information shows the percentage occurrence of the elements in TMs.
@@
Figure 3: Heatmap of the difference in the frequency of elements from $\mathrm{G_{1}}$ and $\mathrm{G_{2}}$ given in Eq. (2)
@@
The single-element $\mathit{\varDelta F}$ analysis, in the preceding discussion, has extended our investigation to consider two-element combinations to further validate the predictive capability of our frequency-based approach and to verify the aforementioned analysis of the dataset $\mathrm{T_{1}}$. Table S3 in the supplementary information lists the various pairs of two-element combinations, the total number of materials containing that pair, the number of TMs having that pair, along with their percentage of occurrence. Our empirical analysis of the dataset $\mathrm{T_{1}}$ in Table S3 reveals that a certain combination of element pairs shows a strong correlation with topological phases such as Pd–Sr and C–Nb, Se–U, Ga–Pt, Pd–Si, and Te–U, etc., suggesting that the co-occurrence of specific elements enhances the likelihood of band inversion and non-trivial topology in the materials. The repeated presence of some elements and pairwise elemental interactions offers additional predictive value, especially for binary materials, and reinforces the reliability of $\mathit{\varDelta F}$ as a descriptor when screening for candidate materials. Furthermore, several of the high-frequency combinations exist with transition metals, heavy pblockp-block elements, alkaline earth metals, etc., which typically have high atomic numbers, spin-orbit coupling, and favorable electronegativity profiles. This two-element analysis also supports the filtering strategy implemented in post-prediction of ML models, where elemental electronegativity and their co-occurrence trends are used to refine model outputs. We have increased confidence in our statistical approach by cross-referencing frequently co-occurring element pairs with $\mathit{\varDelta F<0}$ in the selected candidates upon implementing our trained model on $\mathrm{T_{2}}$.
@@ III.3 Classification of unseen data ($\mathrm{T_{2}}$)
In the preceding section, we have trained and tested it on the labelled data available in the dataset $\mathrm{T_{1}}$; now, we have implemented both models on the unseen data, which is assigned in the dataset $\mathrm{T_{2}}$. It should be noted that the dataset $\mathrm{T_{2}}$ contains the binary combinations of the elements only. As per the observations, the existence of a topological phase depends on the negative value of the $\mathit{\varDelta F}$; hence, using Eq. (5), $\mathit{f_{3}}$ is also likely to be negative, which concludes that all the predicted topological materials possess $\mathit{f_{3}<0}$. The non-trivial topological phase using SVM and RF models has been predicted, respectively, in 84 and 86 materials out of a 1613 binary dataset ($\mathrm{T_{2}}$). We have performed the firstprinciplesfirst-principles calculations of 30 (Table IV and Table V) selected materials out of 84 and 86 topologically non-trivial materials predicted by the SVM and RF models, respectively. The firstprinciplesfirst-principles calculations have shown that these materials have topological semimetallic phases (i.e., DSMs, WSMs, and NLSMs). The crystal structures and band structures of the some selected materials predicted by SVM and RF models are shown in Fig. 4, 5 and Fig. 6, 7, respectively. The corresponding k path selected for band structure calculations in the first Brillouin zone are shown in Fig. S3 and Fig. S4. For the band structure analysis of rest of the systems, Fig. S6 and Fig. S7 can be referred in supplementary information.
Table 4: Materials predicted by SVM to be topological materials (TMs)
@@
Figure 4: Crystal structure of SVM datapoints.
Figure 5: Band structures of materials predicted by SVM: (a) YPd3, (b) AlPd2, (c) Al3Pd5, (d) LuPd3, (e) InPd3, (f) BiPd.
Table 5: Materials predicted by RF to be topological materials (TMs)
@@
Figure 6: Crystal structure of RF datapoints.
Figure 7: Band structures of materials classified by RF: (a) Ti2Re, (b) TePt, (c) YMg, (d) MgSc, (e) SbPt, (f) TiAl.
We have also compared the outcome datapoints of the SVM and RF models and there are 22 such materials that are common in the prediction spaces of both models (Table S6). We have analyzed 19 common materials (excluding the radioactive materials) using firstprinciplesfirst-principles calculations and their topological nature is listed in Table VI. Hence, these materials have the highest probability of a topologically non-trivial phase. The crystal structures and the firstprinciplesfirst-principles calculations of some selected materials from table 6 are shown in Fig. 8 and 9, whereas their k path for band structure calculations are included in Fig. S5. Fig. S8 in the supplementary information can be refereed for the analysis of the rest of the commonly predicted systems.
Table 6: Materials commonly predicted by SVM and RF models to be topological materials (TMs)
@@
Figure 8: Crystal structures of common datapoints.
Figure 9: Band structures of the selected commonly predicted materials using both SVM and RF models: (a) GaPd2, (b) AuPd2, (c) TlZn2, (d) BPd6, (e) SbPt7, (f) ScZn12, (g) Te3Pd10, (h) LuZn12, (i) Pd3Pb, (j) TePd.

### v2

@@ Abstract
The machine learning based approaches efficiently solve the goal of searching the best materials candidate for the targeted properties. The search for topological materials using traditional first-principles and symmetry-based methods often require lots of computing power or is limited by the crystalline symmetries. In this study, we present frequency-based statistical descriptors for machine learning-driven topological material’s classification that is independent of crystallographic symmetry of wave functions. This approach predicts the topological nature of a material based on its chemical formula. With a balanced dataset of 3910 materials, we have achieved classification accuracies of 82% with the Support Vector Machine (SVM) model and 83% with the Random Forest (RF) model, where both models have trained on common frequency based features. We have varified the performances of the models using 5fold5-fold cross-validation approach. Further, we have validated the models on a dataset of unseen binary compounds and have efficiently identified 22 common materials using both the models. Next, we implemented the firstprinciplesfirst-principles approach to confirm the the topological nature of these predicted materials and found the topological signatures of Dirac, Weyl, and nodal-line semimetallic phases. Therefore, we have demonstrated that the implications of frequency-based descriptors is a practical and less complex way to find novel topological materials with certain physical post-processing filters. This approach lays the groundwork for scalable, data-driven topological property screening of complex materials.
@@ I INTRODUCTION
Materials discovery and design are essential in materials science research, which involves finding particularly promising new materials, analyzing and optimizing their attributes for various applications. Topological materials (TMs)[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] have become a subject of immense interest in recent years due to their robust edge/surface states protected by different symmetries from local perturbations such as disorder, impurities, magnetic flux, etc. Since the first realization of TMs in the form of the quantum spin Hall effect [13, 14], this field has remarkably expanded to cover a number of classifications, including topological insulators (TIs) [1, 2, 12, 18], topological semimetals (TSMs),[4, 5, 15, 16, 17, 21, 22, 23], topological crystalline insulators (TCIs) [3, 8, 24] etc. Since the inception of this new class of materials, a significant and persistent investigation has revolved around the methods for ascertaining whether a specific crystalline material possesses a unique topological signature or not. The method of addressing this question has primarily relied on foundational theoretical and experimental techniques that allow trial-and-error possibilities [5, 23, 25]. The first-principles computational approach and traditional experimental methods require a significant amount of time and resources [26, 27, 28, 29, 30, 31, 32]. Recently developed symmetry indicators and topological quantum chemistry theories allow the analysis of a large number of TMs with low computational requirements [33, 34, 35], but despite the effectiveness of these theories, they have limitations owing to the symmetries of the crystalline materials [3, 36]. For example, these theories failed to identify the topological nature of TaAs, which is a low-symmetry crystalline material and the first experimentally identified Weyl semimetal [23]. Moreover, the Chern insulators and the materials having Time Reversal Symmetries (TRS), etc.[14, 37, 38, 39] have robust non-trivial topological phases even when crystalline symmetries vanish. Therefore, we require novel approaches that overcome the aforementioned hurdles, highly accurate and easily executable for the exploration of TMs in condensed matter physics. The machine learning (ML) approach has become one of the adopted methods for screening materials with desired properties [40, 41]. This approach is computationally cheaper, data-driven, with high accuracy and precision, making it an indispensable tool in material discovery [35, 42, 43, 44, 45]. Various reports have been published in recent times that have predicted electronic and structural properties using regression and classification ML models based on chemical composition and crystal structures [36, 46, 47, 48, 49]. Recent advancements include graph-based methods such as crystal graph convolution neural networks (CGCNNs), which have achieved significant success in the prediction and identification of topological phases [50]. A machine-learned chemical descriptor has also been introduced by Ma etet $al.$, thus contributing to the advancements and growing implementation of ML in TMs [51]. The frequency of an atom or an element in a given set of data is an important indicator of what type of elements are more feasible to form TMs. In this study, we use a frequency-based formulation to create a descriptor that can predict TMs with acceptable precision and accuracy. This formulation is straightforward, amenable to calculation, and given materials can be categorized with an accuracy of more than 80% in topological trivial or non-trivial classes. This formulation is independent of the crystal symmetries of the systems and can be implemented on all types of materials. We have integrated our study using ML algorithms, $i.e.$, Support Vector Machine (SVM) and Random Forest (RF), followed by firstprinciplesfirst-principles calculations, for the sake of the verification of the predicted materials. This work aims to improve the accuracy and efficiency of ML models to classify materials for their non-trivial topological phases.
@@ II.1 Data preprocessing
These three quantities in Eq. (2) for each respective element are used to generate the three features for our ML models [refer to section 1.1 and Table S1 in supplementary information] [20]. We generate these features by a weighted sum formula, giving three feature quantities $\mathit{f_{1}}$, $\mathit{f_{2}}$, and $\mathit{f_{3}}$ calculated from the quantities $\mathit{F_{x}^{nt}}$, $\mathit{F_{x}^{rt}}$, and $\mathit{\varDelta F}$ respectively.
@@ II.2 Machine Learning Implementation
Figure 1: The workflow of statistical frequency-based feature engineering approach for training and testing of ML models, and verification of predicted materials using firstprinciplesfirst-principles calculations.
@@ II.3 First-principles calculations
The electronic structure calculations based on the density functional theory (DFT) [54] were performed using the projector augmented-wave (PAW) [55] method as implemented in the Vienna ab initio Simulation Package (VASP) [56] code. The standard PAW potentials, preferably suggested in the VASP code, were used for all the elements for these calculations. The exchange and correlation energy were obtained with the generalized gradient approximation (GGA) of Perdew Burke Ernzerhof (PBE) functional [57], with the inclusion of SOC in the self-consistent field calculations [58]. The total energies convergence criterion of $\mathrm{10^{-6}}$ eV was adopted along with a fine gamma-centered k mesh. The plane-wave cut-off energies were set at 1.5 times the maximum energies available in the PAW potentials.
@@ III.1 Training and Testing
Figure 2: (a, b) Metrics of classification showing Confusion matrices and (c, d) Receiver Operating Characteristic (ROC) - Area Under the Curve (AUC) for the models SVM and RF, respectively. (e, f) ROC-AUC curves for the 5fold5-fold cross-validation of SVM and RF models.
@@ III.1.2 Cross Validation
Cross-validation is a key part of ML analysis, as it permits the removal of reliance on how the sample is divided into testing and training datasets for performance estimation. The cross-validation helps to discover and fix the problem of overfitting and underfitting, and it is a better way to estimate the generalization error of the training-testing split, as it tests the model on more than one validation dataset. This method is used to fine-tune the hyperparameters so they can pick the ones that work best for generalization. Cross-validation also helps make fair comparisons between different models by testing each one on several subsets of data under the same conditions. The complete dataset, consisting of $n$ elements, is randomly divided into $p$ subsets, each containing $n/p$ elements, referred to as folds, as shown in Table III. Here, we have carefully tested the performance of the SVM and RF models by using $\mathit{five-fold}$ cross-validation (FIG. 2 (e, f)), and the AUC scores of each fold, as well as the mean AUC score, are shown in Table III.
@@ III.2 Analysis of data in $\mathrm{T_{1}}$
To support our frequency-based feature approach, we have analyzed the elemental data in the dataset $\mathrm{T_{1}}$ and calculated the elemental features $\mathit{f_{1}}$, $\mathit{f_{2}}$ using Eqs. (3) and (4), whereas the frequency difference $\mathit{\varDelta F}$ and feature $\mathit{f_{3}}$ have been calculated using Eqs. (2) and (5), respectively. We have observed that the $\mathit{\varDelta F}$ values are positive for the alkali metals group (with the exception of hydrogen), whereas the halogens have highly positive $\mathit{\varDelta F}$ values in the pblockp-block elements. Moreover, the highly electronegative elements, such as Oxygen, have an unusually high value of $\mathit{\varDelta F}$, which reflects their widespread occurrence across a varied range of materials in the dataset. The elementary distribution of the rest of the data for $\mathit{\varDelta F}$ across the periodic table can be seen in the heatmap shown in Fig. 3. Among the elements with high negative $\mathit{\varDelta F}$ values, Pd, Pt, Ti, Hf, Zr, U, Au, and Sc occur in more than 50% of the materials containing these elements and are labelled as topologically non-trivial. This observation and analysis imply a higher probability that materials containing these elements possess topological phases. Furthermore, the other elements with negative $\mathit{\varDelta F}$ values generally have a 30% to 50% chance of being in a material of a topological phase. Hence, it implies, statistically, that there is indeed a chance to form topological phases of materials when the elements with $\mathit{\varDelta F<0}$ are present in a material. Moreover, the elements with $\mathit{\varDelta F>0}$ have shown a less than 30% chance of contributing towards this phase. These conclusions are in line with the observation made in the dataset, where most elements with ionic characters are less likely to form topological phases due to their high electronegativity values. This reinforces the significance of the $\mathit{\varDelta F}$ metric as a discriminative feature in identifying potential topological materials and supports its integration into our ML framework. Table S2 in the supplementary information [20] shows the percentage occurrence of the elements in TMs.
@@
Figure 3: Heatmap of the difference in the frequency of elements from $\mathrm{G_{1}}$ and $\mathrm{G_{2}}$ given in Eq. (2).
@@
The single-element $\mathit{\varDelta F}$ analysis, in the preceding discussion, has extended our investigation to consider two-element combinations to further validate the predictive capability of our frequency-based approach and to verify the aforementioned analysis of the dataset $\mathrm{T_{1}}$. Table S3 in the supplementary information [20] lists the various pairs of two-element combinations, the total number of materials containing that pair, the number of TMs having that pair, along with their percentage of occurrence. Our empirical analysis of the dataset $\mathrm{T_{1}}$ in Table S3 reveals that a certain combination of element pairs shows a strong correlation with topological phases such as Pd–Sr and C–Nb, Se–U, Ga–Pt, Pd–Si, and Te–U, etc., suggesting that the co-occurrence of specific elements enhances the likelihood of band inversion and non-trivial topology in the materials. The repeated presence of some elements and pairwise elemental interactions offers additional predictive value, especially for binary materials, and reinforces the reliability of $\mathit{\varDelta F}$ as a descriptor when screening for candidate materials. Furthermore, several of the high-frequency combinations exist with transition metals, heavy pblockp-block elements, alkaline earth metals, etc., which typically have high atomic numbers, spin-orbit coupling, and favorable electronegativity profiles. This two-element analysis also supports the filtering strategy implemented in post-prediction of ML models, where elemental electronegativity and their co-occurrence trends are used to refine model outputs. We have increased confidence in our statistical approach by cross-referencing frequently co-occurring element pairs with $\mathit{\varDelta F<0}$ in the selected candidates upon implementing our trained model on $\mathrm{T_{2}}$.
@@ III.3 Classification of unseen data ($\mathrm{T_{2}}$)
In the preceding section, we have trained and tested it on the labelled data available in the dataset $\mathrm{T_{1}}$; now, we have implemented both models on the unseen data, which is assigned in the dataset $\mathrm{T_{2}}$. It should be noted that the dataset $\mathrm{T_{2}}$ contains the binary combinations of the elements only. As per the observations, the existence of a topological phase depends on the negative value of the $\mathit{\varDelta F}$; hence, using Eq. (5), $\mathit{f_{3}}$ is also likely to be negative, which concludes that all the predicted topological materials possess $\mathit{f_{3}<0}$. The non-trivial topological phase using SVM and RF models has been predicted, respectively, in 84 and 86 materials out of a 1613 binary dataset ($\mathrm{T_{2}}$). We have performed the firstprinciplesfirst-principles calculations of 30 (Table IV and Table V) selected materials out of 84 and 86 topologically non-trivial materials predicted by the SVM and RF models, respectively. The firstprinciplesfirst-principles calculations have shown that these materials have topological semimetallic phases (i.e., DSMs, WSMs, and NLSMs). The crystal structures and band structures of the some selected materials predicted by SVM and RF models are shown in Fig. 4, 5 and Fig. 6, 7, respectively. The corresponding k path selected for band structure calculations in the first Brillouin zone are shown in Fig. S3 and Fig. S4. For the band structure analysis of rest of the systems, Fig. S6 and Fig. S7 can be referred in supplementary information[20].
Table 4: The identified topological phases of the predicted materials with SVM model using firstprinciplesfirst-principles calculations.
@@
Figure 4: The crystal structures of some selected materials from the predicted SVM datapoints.
Figure 5: The projected band structures of some selected materials , $i.e.$, (a) YPd3, (b) AlPd2, (c) Al3Pd5, (d) LuPd3, (e) InPd3, (f) BiPd, from the predicted SVM datapoints.
Table 5: The identified topological phases of the predicted materials with RF model using firstprinciplesfirst-principles calculations.
@@
Figure 6: The crystal structures of some selected materials from the predicted RF datapoints.
Figure 7: The projected band structures of some selected materials , $i.e.$, (a) Ti2Re, (b) TePt, (c) YMg, (d) MgSc, (e) SbPt, (f) TiAl, from the predicted RF datapoints.
We have also compared the outcome datapoints of the SVM and RF models and there are 22 such materials that are common in the prediction spaces of both models (Table S6). We have analyzed 19 common materials (excluding the radioactive materials) using firstprinciplesfirst-principles calculations and their topological nature is listed in Table VI. Hence, these materials have the highest probability of a topologically non-trivial phase. The crystal structures and the firstprinciplesfirst-principles calculations of some selected materials from table 6 are shown in Fig. 8 and 9, whereas their k path for band structure calculations are included in Fig. S5. Fig. S8 in the supplementary information can be refereed for the analysis of the rest of the commonly predicted systems[20].
Table 6: The identified topological phases of the commonly predicted materials with SVM and RF models using firstprinciplesfirst-principles calculations.
@@
Figure 8: The crystal structures of some selected materials from the commonly predicted SVM and RF datapoints.
Figure 9: The projected band structures of some selected materials , $i.e.$, (a) GaPd2, (b) AuPd2, (c) TlZn2, (d) BPd6, (e) SbPt7, (f) ScZn12, (g) Te3Pd10, (h) LuZn12, (i) Pd3Pb, (j) TePd, from the commonly predicted SVM and RF datapoints.